# Complexity Matters: Dynamics of Feature Learning in the Presence of Spurious Correlations

### v1

@@ Abstract
Existing research often posits spurious features as easier to learn than core features in neural network optimization, but the impact of their relative simplicity remains under-explored. Moreover they mainly focus on the end performance intead of the learning dynamics of feature learning. In this paper, we propose a theoretical framework and associated synthetic dataset [^1] The code is made public at:https://github.com/NayutaQiu/Boolean_Spurious grounded in boolean function analysis which allows for fine-grained control on the relative complexity (compared to core features) and correlation strength (with respect to the label) of spurious features to study the dynamics of feature learning under spurious correlation. Our setup uncovers several interesting phenomenon: (1) stronger spurious correlations or simpler spurious features slow down the rate of learning for the core features, (2) learning phases of spurious features and core features are not always separable, (3) spurious features are not forgotten even after core features are fully learned. We show that our findings justify the success of retraining the last layer to remove spurious correlation and also identifies limitations of popular debiasing algorithms that exploit early learning of spurious features. We support our empirical findings with theoretical analyses for the case of learning XOR features with a one-hidden-layer ReLU network.
@@ 1 Introduction
  • Easier spurious features lead to slower core feature emergence. We find that the presence of spurious features notably harm the convergence rate of core feature learning. Spurious features that are easier to learn result in a slower convergence rate compared to their more complex counterparts. This is particularly evident in scenarios where two parity functions of differing degrees are being learned; we give a concrete formula that quantify the initial gradient gap between them. Notably, we found even spurious features that are of similar or slightly lower complexity than the core feature can substantially slow down the convergence rate. Such a delay in convergence can significantly degrade performance, especially in situations where data availability is limited.
  • Core feature learning may not be separable from spurious feature learning. We discover that for staircase functions — a category of functions characterized by their hierarchical structure and learning curves akin to those in real datasets — both core and spurious features are learned concurrently. The degree to which core and spurious features are learned is influenced both by their relative complexity and correlation strength. This observation challenges the effectiveness of widely adopted machine learning algorithms [23, 27, 44, 31, 51] that heavily depend on early learning of shortcut features and make an implicit assumption of a clear-cut separation between the learning phases of core and spurious features.
  • Spurious features are retained. We observe that networks retain spurious features in their representations, even after the core feature has been learned sufficiently well. This retention is particularly notable for spurious features with lower complexity compared to the core features. Not only do these spurious features persist in the network’s representation, but their corresponding weights in the last layer also remain stable, especially under high confounder strength. To provide deeper insight into this issue, we have formulated a general equation that captures the dynamics of core and spurious feature learning in an idealized scenarios.
  • Last Layer Retraining(LLR) decreases reliance on spurious subnetwork: [19, 15] show LLR with balanced dataset is able to improve robustness of the model but the mechanism behind this remains elusive. We demonstrate across numerous datasets that this improvement primarily stems from a reduction in the weights of the last layer that are connected to the spurious subnetwork. Subsequently, we find that applying LLR with heavier $\ell_{1}$ regularization on the original, spurious dataset can also lead to a significant reduction in the influence of the spurious subnetwork. Furthermore, we theoretically quantify the improvement in model robustness through the addition of balanced data.
@@ 1.1 Related Work
Learning under spurious correlation can be interpreted as an Out-Of-Distribution (OOD) or group imbalance task, as spurious features divide the dataset into imbalanced groups. Two cases arise: (1) when the spurious attribute is given, popular methods like [39, 14] can be applied, (2) when the spurious label is unknown during training, various algorithms have been proposed to exploit the phenomenon of simplicity bias [45, 42, 32], which posits that spurious features are learned by the model in the early stages of learning, to upweight underrepresented groups. A representative method of this type is the “Just Train Twice Algorithm”[23], where a model is first trained to identify sample from the minority group.It is worth noting that almost all algorithms assume a balanced validation dataset for extensive hyperparameter tuning, as observed in [15]. Another line of work focuses on underspecified tasks where the spurious features are fully correlated with the label [43, 26].
Last Layer Retraining (LLR).
A key line of work related to our research is [19, 15], where it is demonstrated that LLR on a biased model with balanced data is enough for achieveing state-of-art result on many benmark datasets. [25] further shows this is even true for some benchmark dataset when the spurious data is used. The method essentially runs by first finetune the model then apply logistic regression with heavy regularization term to reweight the last layer. It remains unclear whether the enhanced robustness in previously proposed debiasing algorithms stems from genuine improvement in invariant feature learning or from optimization of the classifier head. Retraining the last layer has also been explored widely and shown to be highly efficient in other settings, such as long-tail learning [22], probing inner representations of a model [1], and out-of-distribution learning [38]. In our study, we assess the quality of the learned representations for both the spurious and core features by evaluating the model’s performance after reweighting following [13].
Learning Boolean functions with Neural Networks (NNs).
@@ 2 Boolean Spurious Features Dataset
$\underbrace{f_{c}:\{+1,-1\}^{c}\to\{+1,-1\}}_{\text{core feature}}\text{ and }\underbrace{f_{s}:\{+1,-1\}^{s}\to\{+1,-1\}}_{\text{spurious feature}}.$
@@
We assume that for both $f\in\{f_{s},f_{c}\}$, the probabilities $P_{U}(f(x)=0)>0$ and $P_{U}(f(x)=1)>0$ (to ensure that the conditional probabilities are well-defined) and $\mathbb{E}_{x\sim D_{U}}[f(x)]=0$ (to be unbiased). A sample in our spurious dataset is $(x,y=f_{c}(x_{c}))$ where $x\sim\mathcal{D}_{\lambda}$. Without loss of generality, we assume $\lambda\geq 0.5$, since $\lambda<0.5$ is symmetric with the spurious feature being $-f_{s}(x_{s})$. In simple words, we can view $\mathcal{D}_{\text{same}}$ as upsampling the distribution that has $f_{s}(x)=f_{c}(x)$ to add spurious correlations. We show some important properties of this distribution in lemma A.1. It is noteworthy that our proposed framework/dataset also satisfies the five constraints proposed in [30] to be considered as a “easy-to-learn” OOD task [^2] The second criterion Identical Invariant Distribution made in [30] may break when $f_{s}(x_{s})$ is biased..
@@
Numerous works have proposed different theoretical and experimental setups to study spurious correlation. It is noteworthy that spurious features are often used interchangeably with ‘shortcut’ or ‘easier’ features. However, different works have drastically different notions to encapsulate the easiness of a feature. For example, [42] examines features along different dimensions, quantifying simplicity by the number of linear segments needed for perfectly separating the data. [46, 50, 40] encapsulate both spurious and core features as 1-bit vectors, gauging simplicity by the amount or variance of noise applied to each feature. Despite our framework bearing resemblance to previously proposed notions of simplicity, we distinguish ourselves by: (1) employing non-linear features for both spurious and core attributes, (2) providing a more general notion of feature complexity and allowing us to explore functions with different properties (3) providing a modular, lightweight implementation of our dataset. Additionally, our dataset allows us to provide theoretical explanation for numerous observed behaviours in learning dynamics under spurious correlation.
We observe that popular semi-synthetic spurious datasets such as Waterbirds[39], Colorful-MNIST [53], and Domino-Image [42] share characteristic learning dynamics shown on the boolean feature datasets (see fig. 4). Therefore, our dataset serves as a good proxy to evaluate algorithms developed to deal with spurious features. Beyond capturing behaviors of existing dataset, our dataset additionally provides precise control over the complexity and structure of the spurious and core functions, which is under-explored in prior datasets. Varying feature structure and complexity results in different training dynamics under spurious correlations. Understanding these differences rigorously allows us to not only refine our understanding of spurious correlations but also helps us understand how algorithms designed to tackle spurious correlations have varying behaviors for different datasets.
Figure 2: Influence of confounder strength and complexity of spurious correlation on learning of core features. The $y$-axis shows the number of epochs required to reach $0.95$ core correlation. For the complexity plot, 0 degree is a control experiment when the model is trained on $\lambda=0.5$. We see spurious feature with lower complexity or higher spurious correlation slow down core feature learning more.
@@ 3.1 Setting
Here, we provide a comprehensive evaluation of a two-layer [^4] Neural Networks with more layers share the same behavior. See appendix C neural network (width 100) optimized using Batch Stochastic Gradient Descent with the cross entropy loss on the boolean features dataset under the online setting. The exact experimental setup can be found in section C.1. We mainly focus on two metrics to measure feature learning:
Core and spurious correlation.
@@
Correlation between the model and core or spurious feature are measured by $\mathbb{E}_{x\sim\mathcal{D}_{\mathsf{unif}}}[f(x)\mathsf{sgn}(h(x))]$ where $\mathsf{sgn}$ is defined to be the sign function, $h$ is the model, and $f$ is either $f_{s}$ or $f_{c}$. Note that $E[f(x)\mathsf{sgn}(h(x))]=P[f(x)=\mathsf{sgn}(h(x))]-(1-P[\mathsf{sgn}(h(x))])$ which is equivalent to $12\cdot\mathsf{accuracy}(f,h)-$ 1. Unlike the common practice in literature, we choose not to use the worst group accuracy as the spurious correlation is symmetric in our cases and the feature distributions are unbiased. In consequence, learning dynamics on subpopulation e.g. $f_{s}(x)=f_{c}(x)=1$ and $f_{s}(x)=f_{c}(x)=0$ are almost identical. Thus the correlation metrics provide a more straightforward view on the behavior of the model.
Decoded core and spurious correlation [19, 13, 38, 1].
@@
We first retrain the last layer of a model with logistic regression to fit either the spurious or core function using a (limited) feature balanced dataset which is $\mathcal{D}_{\lambda}$ with $\lambda=0.5$ in our setting. Then measure the corresponding correlation as above. The decoded correlation metric is used to capture the extent to which a feature’s representation has been effectively learned by the model.
3.2 Results
@@
We summarize our main findings below and provide more detailed experiments report for each finding and claim we make in appendix C. Although our primary focus is on the spurious boolean dataset, we emphasize that each of our findings closely align with observations from other semi-synthetic datasets such as Waterbirds [39], Color-MNIST [53], and different constructions of the Domino Image [42].
@@
(R1) Simpler spurious features and higher correlation strength slow down the convergence rate of core feature learning (fig. 2).
@@
We observe a concave shape relationship between the complexity of spurious features and convergence time, where lower complexity features slow down convergence more significantly. Even when the spurious feature approaches the complexity of the core feature, the model’s performance is still adversely affected by its presence. We also find that slower convergence in learning the core feature leads to poorer overall performance on limited-size datasets. This suggests that the existence of a spurious feature and the strength of the confounder impact the sample complexity required for learning the core feature.
@@
Our investigations indicate that the learning process remains relatively insensitive to the confounder strength until a certain threshold is reached. Beyond this point, there is a substantial increase in the computational time required to learn the core feature. We hypothesize that this threshold phenomenon can be attributed to two factors. Firstly, different features possess varying learning signal strengths as has been hypothesised in various works [37, 35, 40, 32]. Take parity function as an example, differences in gradient signals for features with different complexities are noticeable from initialization. The gradient of a spurious feature can only surpass that of the core feature if the spurious correlation exceeds a certain threshold value. Secondly, as we will explore later, when $\lambda$ is high, it becomes significantly more challenging for gradient descent to either learn the core feature or "unlearn" the spurious feature that has been learned [30].
Figure 3: Weights on spurious and core coordinates in “spurious” and “core” neurons throughout training. Spurious neuron remain spurious while core neurons eventually emerge. Staircase: $deg(f_{s})=7,deg(f_{c})=14,\lambda=0.9$. Parity: $deg(f_{s})=4,deg(f_{c})=10,\lambda=0.9$
(R2) Spurious and core features are learned by two separate sub-networks (fig. 3).
@@
In our setting, spurious feature and core feature learning are disentangled inherently [5]. There exists a classification of neurons into two groups, “spurious neurons” which have larger weights on the spurious index and “core neurons” which have larger weights on the core index in the late stage of learning. The remaining neurons have small weights on all coordinates throughout training. For both parity and staircase tasks, almost all spurious neurons remain focused on spurious coordinates, while core neurons, at the start, do not focus on spurious coordinates and gradually develop an emphasis on core coordinates. The same phenomenon is observed for multi-layer neural network.
For vision task (section C.3.1), it becomes more elusive what neurons are spurious or core. In this case, we retrain the last layer of the neural network on learning either spurious or core feature to separate them apart [15]. We see that the number of set of neurons that has significant weight in both trials is indeed very small and removing them does not hurt the performance of the model in predicting either feature, which suggest that the neurons are separated into two separate networks as in the spurious boolean case.
Figure 4: Core/spurious correlation and decoded correlation dynamics of different datasets. Leftmost figure shows the fourier coefficients of both the spurious and core function are fitted from low (light color) to high (deep color) for the staircase function. The remaining plots demonstrate the improvements achieved through Last Layer Retraining during trainin.And how spurious feature is learned and then memorized. All of the experiments have $\lambda=0.9$. Staircase: $\deg(f_{s})=7,\deg(f_{c})=14$; Parity: $\deg(f_{s})=4,\deg(f_{c})=10$; CIFAR-MINST: (c): Truck-car (s): 01.
(R3) Spurious correlation of the model initially increases and then decreases once core features catch up. However, spurious features remain memorized (fig. 4).
When $\lambda$ is high, the decoded spurious correlation and the total weight within the spurious subnetwork remains high even after extended training, both in the hidden and last layers. This phenomenon persists even when regularization is applied to the loss function  [^5] We note that when regularization term is higher than some threshold value, the core feature learning is infeasible.. Consequently, the learned spurious features continuously influence the learning process of core features. Under the influence, or "slowing down" of the spurious network, the core feature is gradually learned. Eventually, the weights on the core subnetwork surpass those of the spurious network. It’s worth noting that when $\lambda$ is low, around 0.6 in our cases, the decoded spurious correlation and total weights on the spurious subnetwork do decrease significantly after the core feature is learned. In both cases, the learning process for spurious features cease with the increase of core correlation. We will study the learning dynamics of a learned spurious feature later.
@@
We find that last layer retraining consistently improves the worst group accuracy or core function correlation, with the most significant performance boost occurring during the early stages of training (fig. 4). This improvement is attributed to a substantial decrease in the ratio of second-layer weights between spurious neurons and core neurons (table 5). Notably, our findings align with those in [25], where even retraining the last layer on the training dataset (with heavy $\ell_{1}$ regularization) significantly enhances performance. Furthermore, the performance improvement steadily increases when we include more data points from the minority group ($\mathcal{D}_{\text{diff}}$) in the retraining dataset (appenidx fig. 23).
(R5) Popular debiasing algorithms fail in more complex settings.
@@
In scenarios where a spurious attribute is unknown, debiasing algorithms typically adopt a two-stage approach [23, 50, 31, 20, 18] to remove the spurious correlations. They first train an initial model using Early Stop Empirical Risk Minimization (ERM) and then apply different heuristics to distinguish and separate minority group data based on insights from the trained model. They then upweight or upsample the identified minority group data and train a second model. Our dataset uncovers a previously unexplored failure mode of these methods.
These algorithms leverage a notable bias in widely adopted synthetic spurious datasets including Waterbirds [39], Color-MNIST [53], and Domino Images [42], where spurious features are trivial to learn when compared to core features. This leads to a distinct learning process for spurious and core features, demonstrated when the model becomes fully aligned with the spurious feature, as seen in the Domino and Parity experiments (fig. 4). At this stage, misclassified samples correspond precisely to the minority group. Algorithms similar to JTT [23] which upweight misclassified points in the second training phase thus works well under this scenario. However, this distinct phase of learning between spurious and core features generally does not apply, particularly in scenarios with more complex spurious features. An insightful example to illustrate this point is the staircase function where the model fit both of the spurious and core functions from lower degree to higher degree simultaneously (fig. 4). In such cases, unlike parity or cases with trivial spurious feature, misclassified points may not accurately represent the minority group.
In order to make this argument rigorous, we evaluate the quality of the inferred minority group using Jaccard scores defined as $\frac{|A\cap B|}{|A\cup B|}$ and containment scores $\frac{|A\cap B|}{|A|}$, where $A$ represents the predicted minority group by the algorithm using an early stop model, and $B$ represents the ground truth minority group to evaluate their performance with model stopped at each training epoch on different datasets (see fig. 5). We observe that various debiasing algorithms [23, 25, 50] do not attain scores above 0.5 throughout training on the staircase setup. Upweighting the inferred minority group therefore results in over half of the samples from the majority group being upweighted. This can potentially introduce unknown biases into the model and result in the failure of core feature learning. Contrast this to the parity or easy spurious feature case where the scores reach almost 1.
Building on these insights, we design a more challenging domino-vision dataset which features a more complex spurious task and demonstrate that the model fails to show further improvement when applied with current debiasing algorithms [25, 23, 50]. This suggests that our dataset serves as a good benchmark for future debiasing algorithms to explore scenarios with more complex spurious features. For a more detailed discussion of the limitations of previous debiasing algorithms and their performance on other spurious datasets, see section D.1.
@@
Figure 5: The two plots on the right demonstrate that debasing methods JTT [23] and Spare [50] is able to infer points from minority group successfully on datasets that share the characteristic of the parity case where the spurious feature is much easier than the core feature. And there exist a point where the model is fully correlated with the spurious feature. While for the spurious staircase experiment and more challenging Domino dataset shown in the left (taking core feature as cat-dog and spurious feature as truck-automobile from the CIFAR dataset [21]), the highest Jaccard score remains below 0.5. Here the black dash line remark the minority group proportion in the training dataset.

### v2

@@ Abstract
Existing research often posits spurious features as easier to learn than core features in neural network optimization, but the impact of their relative simplicity remains under-explored. Moreover, studies mainly focus on end performance rather than the learning dynamics of feature learning. In this paper, we propose a theoretical framework and an associated synthetic dataset [^1] The code is made public at:https://github.com/NayutaQiu/Boolean_Spurious grounded in boolean function analysis. This setup allows for fine-grained control over the relative complexity (compared to core features) and correlation strength (with respect to the label) of spurious features to study the dynamics of feature learning under spurious correlations. Our findings uncover several interesting phenomena: (1) stronger spurious correlations or simpler spurious features slow down the learning rate of the core features, (2) two distinct subnetworks are formed to learn core and spurious features separately, (3) learning phases of spurious and core features are not always separable, (4) spurious features are not forgotten even after core features are fully learned. We demonstrate that our findings justify the success of retraining the last layer to remove spurious correlation and also identifies limitations of popular debiasing algorithms that exploit early learning of spurious features. We support our empirical findings with theoretical analyses for the case of learning XOR features with a one-hidden-layer ReLU network.
@@ 1 Introduction
  • Easier spurious features lead to slower core feature emergence. We find that the presence of spurious features notably harm the convergence rate of core feature learning when infinite data is available. This is particularly evident in scenarios where two parity functions of differing degrees are being learned; we give a concrete formula that quantify the initial gradient gap between them. Notably, we found even spurious features that are of similar or slightly lower complexity than the core feature can substantially slow down the convergence rate. This delay in convergence manifests as poor robustness of the model when data availability is limited.
  • The common assumption that the learning phase is separated into spurious feature learning and then core feature learning can lead to complete failure of various debiasing algorithm. [26, 30, 46, 34, 53] heavily depend on early learning of shortcut features and make an implicit assumption of a clear-cut separation between the learning phases of core and spurious features. We show that this assumption is generally incorrect and provide an insightful counterexample to demonstrate how such algorithms can fail completely. Staircase functions—a category of functions characterized by their hierarchical structure and learning curves similar to those in real datasets—illustrate that both core and spurious features are learned concurrently. The degree to which core and spurious features are learned is influenced by their relative complexity and correlation strength. This observation challenges the effectiveness of widely adopted machine learning algorithms.
  • Spurious features are retained. We observe that networks retain spurious features in their representations, even after the core feature has been learned sufficiently well. This retention is particularly notable for spurious features with lower complexity compared to the core features. Not only do these spurious features persist in the network’s representation, but their corresponding weights in the last layer also remain stable, especially under high confounder strength. We provide theoretical explanation for such phenomenon and show how well the spurious feature is being memorized is closely related to the correlation strength.
  • The network is separated into two distinct subnetworks in learning different features, and Last Layer Retraining (LLR) decreases reliance on the spurious subnetwork: [20, 16] show LLR with balanced dataset is able to improve robustness of the model. While it is clear from the previous works that the core feature can be linearly decoded from the last layer, the mechanism behind this remains elusive. We demonstrate across numerous datasets that this improvement primarily stems from a reduction in the weights of the last layer that are connected to the spurious subnetwork. This observation is based on the finding that spurious and core representations are disentangled in the last layer.
@@ 1.1 Related Work
Learning under spurious correlation can be interpreted as an Out-Of-Distribution (OOD) or group imbalance task, as spurious features divide the dataset into imbalanced groups. Two cases arise: (1) when the spurious attribute is given, popular methods like [41, 15] can be applied, (2) when the spurious label is unknown during training, various algorithms have been proposed to exploit the phenomenon of simplicity bias [47, 44, 35], which posits that spurious features are learned by the model in the early stages of learning, to upweight underrepresented groups. A representative method of this type is the “Just Train Twice Algorithm”[26], where a model is first trained to upweight “easy” samples.It is worth noting that almost all algorithms assume a balanced validation dataset for extensive hyperparameter tuning, as observed in [16]. Another line of work focuses on underspecified tasks where the spurious features are fully correlated with the label and attempt to learn diverse features [45, 29].
Last Layer Retraining.
A key line of work related to our research is [20, 16], where it is demonstrated that last layer retraining on a biased model with balanced data is enough for achieveing state-of-art result on many benmark datasets. [28] further shows this is even true for some benchmark dataset when the spurious data is used. The method essentially runs by first finetune the model then apply logistic regression on a group balanced dataset with heavy regularization term to reweight the last layer. Retraining the last layer has also been explored widely and shown to be highly efficient in other settings, such as long-tail learning [24], probing inner representations of a model [1], and out-of-distribution learning [40]. In our study, we assess the quality of the learned representations for both the spurious and core features by evaluating the model’s performance after reweighting following [14].
Learning Boolean functions with Neural Network(NN).
@@ 2 Boolean Spurious Features Dataset
$\underbrace{f_{c}:\{+1,-1\}^{c}\to\{+1,-1\}}_{\text{core feature}}$
$\underbrace{f_{s}:\{+1,-1\}^{s}\to\{+1,-1\}}_{\text{spurious feature}}$
@@
We assume that for both $f\in\{f_{s},f_{c}\}$, the probabilities $P_{U}(f(x)=0)>0$ and $P_{U}(f(x)=1)>0$ (to ensure that the conditional probabilities are well-defined) and $\mathbb{E}_{x\sim D_{U}}[f(x)]=0$ (to be unbiased). A sample in our spurious dataset is $(x,y=f_{c}(x_{c}))$ where $x\sim\mathcal{D}_{\lambda}$. Without loss of generality, we assume $\lambda\geq 0.5$, since $\lambda<0.5$ is symmetric with the spurious feature being $-f_{s}(x_{s})$. In simple words, we can view $\mathcal{D}_{\text{same}}$ as upsampling the distribution that has $f_{s}(x)=f_{c}(x)$ to add spurious correlations. It is helpful to recognize that $x_{s}-f(x_{s})-y=f(x_{c})-x_{c}$ forms a markov chain. We show some important properties of this distribution in lemma A.1. It is noteworthy that our proposed framework/dataset also satisfies the five constraints proposed in [33] to be considered as a “easy-to-learn” OOD task [^2] The second criterion Identical Invariant Distribution made in [33] may break when $f_{s}(x_{s})$ is biased..
@@
Figure 2: Core/spurious correlation and decoded correlation dynamics of different datasets. Leftmost figure shows the fourier coefficients of both the spurious and core function are fitted from low (light color) to high (deep color) for the staircase function. All of the experiments have $\lambda=0.9$. Staircase: $\deg(f_{s})=7,\deg(f_{c})=14$; Parity: $\deg(f_{s})=4,\deg(f_{c})=10$; CIFAR-MINST: (c) Truck-car (s) 01.
@@
Numerous works have proposed different theoretical and experimental setups to study spurious correlation. It is noteworthy that spurious features are often used interchangeably with ‘shortcut’ or ‘easier’ features. However, different works have drastically different notions to encapsulate the easiness of a feature. For example, [44] examines features along different dimensions, quantifying simplicity by the number of linear segments needed for perfectly separating the data. [48, 52, 42, 6] encapsulate both spurious and core features as 1-bit vectors, gauging simplicity by the amount or variance of noise applied to each feature. Despite our framework bearing resemblance to previously proposed notions of simplicity, we distinguish ourselves by: (1) employing non-linear features for both spurious and core attributes, (2) providing a more general notion of feature complexity and allowing us to explore functions with different properties (3) providing a modular, lightweight implementation of our dataset. Additionally, our dataset allows us to provide theoretical explanation for numerous observed behaviours in learning dynamics under spurious correlation.
We observe that popular semi-synthetic spurious datasets such as Waterbirds[41], Colorful-MNIST [56], and Domino-Image [44] share characteristic learning dynamics shown on the boolean feature datasets (see fig. 2). Therefore, our dataset serves as a good proxy to evaluate algorithms developed to deal with spurious features. Beyond capturing behaviors of existing dataset, our dataset additionally provides precise control over the complexity and structure of the spurious and core functions, which is under-explored in prior datasets.
@@ 3 Empirical Findings
Here, we provide a comprehensive evaluation of a two-layer [^4] Neural Networks with more layers share the same behavior. See Figure 9. neural network (width 100) optimized using Batch Stochastic Gradient Descent with the cross entropy loss on the boolean features dataset under the online setting. The exact experimental setup can be found in the Appendix C.1. We emphasize that our main focus here is on the online setting, and we provide more experimental findings regarding limited dataset size in the appendix. We mainly focus on two metrics to measure feature learning:
@@
Core and spurious correlation: correlation between the model and core or spurious feature are measured by $\mathbb{E}_{x\sim\mathcal{D}_{\mathsf{unif}}}[f(x)\mathsf{sgn}(h(x))]$ where $\mathsf{sgn}$ is defined to be the sign function, $h$ is the model, and $f$ is either $f_{s}$ or $f_{c}$. Note $\mathcal{D}_{\mathsf{unif}}$ is a group balanced distribution as the functions we studied are unbiased. Thus the core correlation is exactly the mean group accuracy, a metric extensively used in the literature. Additionally, since the spurious correlation is symmetric in our setting, the core correlation closely matches the worst group accuracy.
@@
Decoded core and spurious correlation [20, 14, 40, 1]: we first retrain the last layer of a model with logistic regression to fit either the spurious or core function using a group balanced dataset. Then measure the corresponding correlation as above. The decoded correlation metric is used to capture the extent to which a feature’s representation has been effectively learned by the model.
@@
Although our primary focus is on the spurious Boolean dataset, we emphasize that our findings closely align with observations from other semi-synthetic datasets such as Waterbirds [41], CMNIST[56], and Domino Image[44]. For more detailed exploration of these datasets, see appendix C.
Figure 3: Influence of confounder strength and complexity of spurious correlation on learning of core features. The $y$-axis shows the number of epochs required to reach $0.95$ core correlation. The 00 degree bar indicate the epochs required to learn core feature when spurious correlation is not present i.e $\lambda=0.5$. Each bar of the left two plots is based on 30 repetitions of experiments.
@@
(R1) Simpler spurious features and higher correlation strength slow down the convergence rate of core feature learning
@@
(Figure 3). We observe a concave U-shaped phenomenon in the relationship between the complexity of spurious features and convergence time, where lower complexity features slow down convergence. Remarkably, even when the spurious feature approaches the complexity of the core feature, the model’s performance is still adversely affected by its presence. Additionally, we find that slower convergence in learning the core feature leads to poorer overall performance on limited-size datasets (see Table C.2.1). This suggests that the existence of a spurious feature impact the sample complexity required for learning the core feature.
@@
Our investigations indicate that the learning process remains relatively insensitive to the confounder strength until a certain threshold is reached. Beyond this point, there is a sudden and substantial increase in the computational time required to learn the core feature [^5] Note that our experiments show that even when the confounder strength is as high as 0.99, the model eventually fits the core function perfectly. Refer to Appendix C.2.2 for more details. We hypothesize that this threshold phenomenon can be attributed to two factors. Firstly, different features possess varying learning signal strengths. In the simplest case, exemplified by the parity function, differences in gradient signals for features with different complexities are noticeable from initialization. The gradient of a spurious feature can only surpass that of the core feature if the spurious correlation exceeds a certain threshold value. Secondly, as we will explore later, when $\lambda$ is high, it becomes significantly more challenging for gradient descent to "unlearn" spurious neurons.
(R2) Spurious and core features are learned by two separate sub-networks
@@
(Figure 4). There exists a classification of neurons into two groups, “spurious neurons” which have larger weights on the spurious index and “core neurons” which have larger weights on the core index in the late stage of learning. For both parity and staircase tasks, almost all spurious neurons remain focused on spurious coordinates, while core neurons, at the start, do not focus on spurious coordinates and gradually develop an emphasis on core coordinates. See Appendix C.3.1 for more detail.
In vision tasks, it becomes more challenging to identify spurious or core neurons. To address this, we retrain the last layer of the neural network to learn either the spurious or core function separately. We observe that the set of neurons with significant weights in both trials is indeed very small, suggesting that neurons are separated into two distinct networks, similar to the spurious Boolean case. Our studies indicate that non-causally related feature representations are perhaps disentangled (at least in the last layer) from the outset, which aligns with a common goal in the fairness and model explainability literature [25, 13]. Therefore, it is of future interest to understand what conditions are sufficient for a model to learn disentangled representations under common training procedures.
Figure 4: Each plot here shows the weight dynamic throughout training within a single selected neuron and each curve here corresponding to the weight dynamic on a single coordinate. The left two plots are for staircase with $\text{deg}(f_{s})=7,\text{deg}(f_{c})=14$ and the right two plots are for parity with $\text{deg}(f_{s})=4,\text{deg}(f_{c})=10$. For both experiment, $\lambda=0.9$. We see the neurons are separated into core and spurious neurons. Spurious neurons remain focus on learning spurious feature and core neuron eventually emerge and learns the core feature.
(R3) Spurious correlation strength determine how well the spurious feature is memorized.
(fig. 2, fig. 5). When $\lambda$ is high, the decoded spurious correlation and the total weight within the spurious subnetwork remains high even after extended training, both in the hidden and last layers. This phenomenon persists even when regularization is applied to the loss function.
Notably, when $\lambda$ is low, around 0.6 in our cases, the decoded spurious correlation and total weights on the spurious subnetwork decrease over time as the core feature is learned. In both cases, the learning process for spurious features plateaus when the core correlation starts to exceed the spurious correlation. Our observation thus illustrates another kind of in-distribution forgetting that occurs during training, contrasting with the established catastrophic forgetting, which happens when training on out-of-distribution (OOD) data. Thus to learn diverse features, it could be beneficial to identify and freeze such spurious neurons adaptively as have been done in [22, 54].
Figure 5: The two plots are produced by running experiments on parity cases under different $\lambda$, the focus here is on the decoded spurious correlation. The experiments are run under $\text{deg}(f_{s})=4,\text{deg}(f_{c})=10$. Left: we see when $\lambda$ is low, the spurious feature is being forgotten in the later stage of training. Right: when $\lambda$ is relatively high, the spurious feature is memorized once it is learned.
@@
We observed that last layer retraining consistently improves the worst group accuracy or core function correlation, with the most significant performance boost occurring during the early stages of training (Figure 2). This improvement is attributed to a substantial decrease in the ratio of second-layer weights between spurious neurons and core neurons (Table 5) which is a consequence of R2 and R3.
Notably, our findings align with those in [28], where we observed that even retraining the last layer on the training dataset (with heavy $l1$ regularization) significantly enhances robustness. Furthermore, we found the performance boost is most significant when just a small amount of group-balanced data is used for LLR (fig. 24).
(R5) Popular debiasing algorithms fail in more general settings.
@@
(Figure 6) In scenarios where a spurious attribute is absent, debiasing algorithms typically adopt a two-stage approach [26, 52, 34, 21, 19]. They first train an initial model using Early Stop Empirical Risk Minimization (ERM). These algorithms diverge in the second stage, where different heuristics are applied to distinguish and separate minority group data based on insights from the initial model. Implicit in their approach is the assumption of an extreme bias toward simplicity in the spurious feature, expecting the early model to prioritize learning the spurious feature and providing valuable information for segregating minority group samples.
However, our investigation reveals that the benchmark datasets commonly used by these algorithms exhibit a crucial dataset bias towards having a much simpler spurious feature compared to the core feature. This bias creates a distinct separation between the learning phases of spurious and core features, as demonstrated in our parity case, allowing these methods to effectively separate minority groups (see Figure 6). We demonstrate that this separation may not hold true in many cases, particularly with limited datasets and spurious features of similar complexity to the core feature, as illustrated in the staircase case. To emphasize the practicality of our dataset, we introduce a domino-vision dataset with more challenging spurious and core features and shows the model is not able to improve further using early stopped model. To further assess the effectiveness of these debiasing algorithms, we employ Jaccard scores defined as $\frac{|A\cap B|}{|A\cup B|}$ and containment scores $\frac{|A\cap B|}{|A|}$, where $A$ represents the predicted minority group by the algorithm using an early stop model, and $B$ represents the ground truth minority group. These metrics allow us to evaluate the extent to which minority group data is included in the predictions. For a more detailed discussion of the weaknesses of previous debiasing algorithms and their performance on real datasets, please refer to Appendix D.1.
@@
Figure 6: The two plots on the right demonstrate that debasing methods JTT [26] and Spare [52] is able to infer points from the minority group successfully on datasets that share the characteristic of the parity case where there is an early spike in spurious correlation and the spurious feature is much easier than the core feature. While for the spurious staircase experiment and more challenging Domino dataset shown in the left the highest Jaccard score remains below 0.5. We report the worst group accuracy is not improving after upsampling based on the inferred group. See section D.1 for more detail. The black dash line remark the minority group proportion in the training dataset.